OpenStudy (anonymous):

x+y+z=2 y-3z=6 2x-3y+z=6 HELP!!

OpenStudy (anonymous):

there is no easy way to do this other than cheating

OpenStudy (anonymous):

Can I get help step by step?

OpenStudy (uri):

Get all the x values together,all y together and all numericals together and all z together.

OpenStudy (anonymous):

we can try i guess you have to use elimination

OpenStudy (uri):

Also if you're shifting after = <-- sign changes

OpenStudy (anonymous):

alright, I'm not sure how exactly to solve this.. I'll appreciate it!

OpenStudy (anonymous):

If I get help step by step. (:

OpenStudy (anonymous):

\[x+y+z=2\\ y-3z=6 \\2x-3y+z=6\]

OpenStudy (uri):

Is that Cramers rule? @satellite73

OpenStudy (anonymous):

lets get rid of the \(x\) first by taking \[x+y+z=2\\2x-3y+z=6\] and multiply the first equation by \(-2\) to get \[-2x-2y-2z=-4\\2x-3y+z=6\]

OpenStudy (agent0smith):

You can use elimination to get rid of x in the first and third equations. Then you have a simpler system of two equations with y and z only. But there's many ways to do this.

OpenStudy (anonymous):

add \[-2x-2y-2z=-4\\2x-3y+z=6\] and get \[-5y-z=2\]now the \(x\) is gone

OpenStudy (anonymous):

now looking at the second equation and the one above, we get two equations in two unknowns \[-5y-z=2\\y-3z=6\]

OpenStudy (anonymous):

solving this almost in your head you get \(y=0, z=-2\) but we can work out those steps too

OpenStudy (anonymous):

alrighty (:

OpenStudy (anonymous):

\[-5y-z=2\\y-3z=6\] multiply the second one by 5 get \[-5y-z=2\\5y-15z=30\] add and get \(-16z=32\iff z=-2\)

OpenStudy (anonymous):

then substitute back to find \(y-3(-2)=6\implies y+6=6\implies y=0\) as promised

OpenStudy (anonymous):

now you have \(y=0, z=-2\) substitute in to \[x+y+z=2\] and get \[x+0-2=2\implies x=4\]

OpenStudy (anonymous):

solution is \((4,0,-2)\) and i hope i never have to do another one of these as long as i live

OpenStudy (anonymous):

haha thank you so much!

OpenStudy (anonymous):

if it was me, i would do this http://www.wolframalpha.com/input/?i=x%2By%2Bz%3D2%2C+y-3z%3D6%2C+2x-3y%2Bz%3D6+ done is 2 seconds, and have a snack

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thank you again!! i have more to do but i kinda got the idea!

OpenStudy (anonymous):

@satellite73 How do I use that website?